# Visual Hand Gesture Recognition with Deep Learning: A Comprehensive Review of Methods, Datasets, Challenges and Future Research Directions

### v2

@@ 11 Conclusions
Nevertheless, the most important challenges regarding VHGR are not yet solved, leaving substantial research gaps; the data scarcity - especially concerning more specialized tasks, such as the recognition of rare sign languages - and the high computational cost remain essentially unresolved. Furthermore, the existing methods are limited to a predefined list of gestures, which may not be optimal for real-world interaction with humanoid robots or interpretation of the body language. Thus, VHGR could be expected to open-vocabulary settings, enabling deployment to broader applications. The potential of leveraging the extensive knowledge of foundation models (kas2025foundation_models) in the context of VHGR enabling zero-shot gesture recognition as well as the utilization of model compression techniques (DL_MODEL_COMPRESSION_marino2023deep) to reduce the computational complexity are also rarely explored, leaving space for future work.
@@ Acknowledgments
Author contribution Konstantinos Foteinos, Jorgen Cani, Manousos Linardakis, Iraklis Varlamis and Georgios Th. Papadopoulos performed the literature review and prepared a draft of the manuscript. Panagiotis Radoglou-Grammatikis, Vasileios Argyriou and Panagiotis Sarigiannidis performed reviewing and editing of the manuscript. Georgios Th. Papadopoulos was responsible for receiving the funding to implement the study.

### v3

@@ 11 Conclusions
Nevertheless, the most important challenges regarding VHGR are not yet solved, leaving substantial research gaps; the data scarcity - especially concerning more specialized tasks, such as the recognition of rare sign languages - and the high computational cost remain essentially unresolved. Furthermore, the existing methods are limited to a predefined list of gestures, which may not be optimal for real-world interaction with humanoid robots or interpretation of the body language. Thus, VHGR could be expected to open-vocabulary settings, enabling deployment to broader applications. The potential of leveraging the extensive knowledge of foundation models (97) in the context of VHGR enabling zero-shot gesture recognition as well as the utilization of model compression techniques (129) to reduce the computational complexity are also rarely explored, leaving space for future work.
@@ Acknowledgments
Author contribution Konstantinos Foteinos, Manousos Linardakis, Iraklis Varlamis and Georgios Th. Papadopoulos performed the literature review and prepared a draft of the manuscript. Panagiotis Radoglou-Grammatikis, Vasileios Argyriou and Panagiotis Sarigiannidis performed reviewing and editing of the manuscript. Georgios Th. Papadopoulos was responsible for receiving the funding to implement the study.